OpenStudy (17bohtal):

Find the approximate length of the leg of a right triangle with one leg length 8 and hypotenuse length 19.

OpenStudy (geekfromthefutur):

Ok so we got two parts here

OpenStudy (17bohtal):

okie.

OpenStudy (geekfromthefutur):

My bad sorry for the confusion lol i havent had coffee .

OpenStudy (17bohtal):

Its all good haha

OpenStudy (geekfromthefutur):

Hypotenuse = c a2+b2=c282+92=c2=145c=145and then that equals????

OpenStudy (17bohtal):

.... whut...

OpenStudy (phi):

First, write down the pythagorean theorem (very famous) \[ a^2 + b^2 = c^2 \] a and b are the "legs" c is the hypotenuse. replace the letter a with the length of the leg replace the letter c with the len of the hypotenuse can you do that ?

OpenStudy (17bohtal):

Okay I got it.

OpenStudy (geekfromthefutur):

Like this basically a^2+b^2=c^2 8^2+9^2=c^2=145 c=√145=__.__

OpenStudy (17bohtal):

I dont know who I should give the medal to . You both did very well.

OpenStudy (geekfromthefutur):

Well i gave phi a medal so win/win but the medal doesn't really matter

OpenStudy (phi):

*** Okay I got it.*** what did you get. Can you write it down ? replace the letter a with the length of the leg replace the letter c with the len of the hypotenuse

OpenStudy (geekfromthefutur):

Its about who will help you get the answer correct and get it done as a team

OpenStudy (geekfromthefutur):

I got it too

OpenStudy (17bohtal):

I know how to do it now. Thanks guys!

OpenStudy (geekfromthefutur):

What did you get tho?

OpenStudy (phi):

trust me, you should work through this problem step by step.

OpenStudy (17bohtal):

20.61

OpenStudy (geekfromthefutur):

Ur wrong

OpenStudy (17bohtal):

How?

OpenStudy (geekfromthefutur):

its 12.04

OpenStudy (17bohtal):

Thats even a choice for the multiple choice.

OpenStudy (phi):

1) write down the equation a^2 + b^2 = c^2 a and b are the legs, c is the hypotenuse 2) replace a with 8 and c with 19

OpenStudy (geekfromthefutur):

Ik what i did wrong i did it for a 9 not 19 im ashamed listen to phi he/she knows what to do

OpenStudy (17bohtal):

Its 17.2

OpenStudy (17bohtal):

correct?

OpenStudy (phi):

to "replace" a with 8 means erase the a and put in 8 replace c with 19 8^2 + b^2 = 19^2

OpenStudy (phi):

now simplify 8^2 = 8*8 simplify 19^2 = 19*19 what do you get ?